prettydevil45:

The slope of the line whose equation is 3y = 2x - 3 is 2 2/3 3/2

RKY22:

https://www.tiger-algebra.com/ try this site it might help

darkknight:

3y = 2x - 3 is divide both sides by 3 to solve for y y = (2/3)(x)-1 slope = 2/3